# and Ioannis Patras

### v1

@@ IV Experiments
SINet [1] CVPR 20 F 0.092 0.745 0.804 0.644 0.043 0.776 0.864 0.631 0.058 0.808 0.871 0.723
UGTR [41] ICCV 21 F 0.086 0.784 0.822 0.684 0.036 0.817 0.852 0.666 0.052 0.839 0.874 0.747
MGL-R [42] CVPR 21 F 0.088 0.775 0.812 0.673 0.035 0.814 0.851 0.666 0.052 0.833 0.867 0.740
PFNet [43] CVPR 21 F 0.085 0.782 0.841 0.695 0.040 0.800 0.877 0.660 0.053 0.829 0.887 0.745
UJSC [44] CVPR 21 F 0.073 0.800 0.859 0.728 0.035 0.809 0.884 0.684 0.047 0.842 0.898 0.771
BSA-Net [45] AAAI 22 F 0.079 0.794 0.851 0.763 0.034 0.818 0.891 0.738 0.048 0.841 0.897 0.808
ZoomNet [5] CVPR 22 F 0.066 0.820 0.892 0.752 0.029 0.838 0.911 0.729 0.043 0.853 0.896 0.784
SINetv2 [9] TPAMI 22 F 0.070 0.820 0.895 0.782 0.037 0.815 0.906 0.718 0.048 0.847 0.914 0.805
HitNet [46] AAAI 23 F 0.055 0.849 0.906 0.831 0.023 0.871 0.935 0.823 0.037 0.875 0.926 0.853
SAM-Ada. [32] ICCV 23 F 0.070 0.847 0.873 0.765 0.025 0.883 0.918 0.801 - - - -
FEDER [16] CVPR 23 F 0.071 0.802 0.898 0.781 0.032 0.822 0.905 0.751 0.044 0.808 0.871 0.769
FSPNet [17] CVPR 23 F 0.050 0.856 0.899 0.830 0.026 0.851 0.895 0.769 0.035 0.879 0.915 0.843
DINet [14] ECCV 24 F 0.068 0.821 0.874 0.755 0.031 0.832 0.903 0.729 0.043 0.856 0.909 0.794
CamoDiffusion [20] AAAI 24 F 0.042 0.878 0.940 0.853 0.020 0.881 0.944 0.814 0.029 0.893 0.942 0.859
CamoFormer [19] TPAMI 24 F 0.046 0.872 0.929 0.854 0.023 0.869 0.932 0.811 0.030 0.891 0.939 0.868
MCRNet [21] IJCV 25 F 0.041 0.883 0.939 0.849 0.020 0.880 0.941 0.812 0.029 0.895 0.943 0.857
SCSOD [47] AAAI 21 S 0.102 0.713 0.795 0.618 0.055 0.710 0.805 0.546 - - - -
CRNet [10] AAAI 23 S 0.092 0.735 0.815 0.641 0.049 0.733 0.832 0.576 0.063 0.775 0.855 0.688
SAM [13] ICCV 23 - 0.132 0.684 0.687 0.606 0.050 0.783 0.798 0.701 0.078 0.767 0.776 0.696
SAM-S [13] ICCV 23 S 0.105 0.731 0.774 - 0.046 0.772 0.828 - 0.071 0.763 0.832 -
WS-SAM [11] NeurIPS 23 S 0.092 0.759 0.818 0.667 0.038 0.803 0.878 0.680 0.052 0.829 0.886 0.757
SAM-COD [12] ECCV 24 S 0.060 0.836 0.903 0.779 0.029 0.833 0.904 0.728 0.039 0.859 0.912 0.795
@@ IV-A1 Implementation Details
We implement our model using PyTorch [48] and conduct all experiments on a single NVIDIA RTX 3090 GPU. In the first stage, we adopt SAM [13] with a ViT-H backbone and Qwen2.5-VL-7B [49] as the multimodal large language model (MLLM). In the second stage, following previous studies [17, 50], we employ the base version of ViT [51] pretrained with the DeiT strategy [52] as the transformer encoder.
For optimization, we use stochastic gradient descent (SGD) [53] with a momentum of 0.9 and a weight decay of $5\times 10^{-4}$. The batch size is set to 4, and the initial learning rate is 0.03. The learning rate is linearly warmed up during the first 20 epochs and subsequently decayed using a cosine annealing schedule. For loss configurations, the weight $w_{n}$ of the non-scribble class in $\mathcal{L}_{\text{Scrib}}$ is set to 0.02, and the modulation factor $\gamma$ in $\mathcal{L}_{\text{Debias}}$ is 0.9. In the overall loss formulation, the balancing coefficients are set to $\alpha=2$ and $\beta=0.5$.
@@ IV-A2 Datasets
To evaluate the effectiveness of our proposed method, we leverage the scribble-annotated S-COD dataset [10], which is specifically designed for weakly-supervised camouflaged object detection, to generate pseudo masks and train our model. The trained model is then evaluated on three widely used benchmark datasets: CAMO [54], COD10K [9], and NC4K [55]. CAMO consists of 1,000 training images and 250 testing images. COD10K is a larger and more challenging dataset containing 3,040 training images and 2,026 testing images. NC4K serves as an additional benchmark, comprising 4,121 testing images without a training split.
@@ IV-A3 Evaluation Metrics
Following existing works [11, 12, 10], we adopt four commonly used metrics for evaluation: mean absolute error (MAE), structure measure ($S_{m}$) [56], mean E-measure ($E_{m}$) [57], and weighted F-measure ($F_{\beta}^{w}$) [58].
@@ IV-B1 Quantitative Comparison
To verify the effectiveness of our proposed framework, we compare it against state-of-the-art fully supervised and weakly-supervised methods, including SINet [1], UGTR [41], MGL-R [42], PFNet [43], UJSC [44], BSA-Net [45], ZoomNet [5], SINetv2 [9], HitNet [46], SAM-Ada [32], FEDER [16], FSPNet [17], DINet [14], CamoDiffusion [20], CamoFormer [19], MCRNet [21], SCSOD [47], CRNet [10], SAM [13], SAM-S [13], WS-SAM [11], and SAM-COD [12].
From the results in Table I, it is evident that ${D}^{3}$ETOR consistently outperforms all scribble-supervised baselines across all datasets. For example, on CAMO, it achieves an MAE of 0.045, $S_{m}$ of 0.868, $E_{m}$ of 0.928, and $F_{\beta}^{w}$ of 0.823, surpassing the second-best scribble-supervised method SAM-COD [12] (MAE 0.060, $S_{m}$ 0.836, $E_{m}$ 0.903, $F_{\beta}^{w}$ 0.779) by a clear margin.
Furthermore, ${D}^{3}$ETOR demonstrates competitive or superior performance compared to most fully supervised approaches, only slightly behind CamoDiffusion [20] and MCRNet [21] on certain metrics. Overall, these results indicate that ${D}^{3}$ETOR establishes a new state-of-the-art for camouflaged object detection in weakly and even fully supervised scenarios using only scribble annotations.
@@ IV-B2 Qualitative Comparison
As illustrated in Figure 8, we select all existing weakly supervised approaches (i.e., CRNet [10], WS-SAM [11], and SAM-COD [12]) to visually compare against our proposed ${D}^{3}$ETOR. The qualitative results indicate that ${D}^{3}$ETOR accurately segments camouflaged objects under a variety of challenging conditions, including small-scale targets, intricate structural details, high visual resemblance to the background, fuzzy object boundaries, and complex surrounding contexts. In contrast, other methods fail to capture complete object regions or produce predictions with blurred details. These qualitative analyses provide an intuitive demonstration of the effectiveness and robustness of ${D}^{3}$ETOR.

### v2

@@ IV Experiments
SINet [fan2020camouflaged] CVPR 20 F 0.092 0.745 0.804 0.644 0.043 0.776 0.864 0.631 0.058 0.808 0.871 0.723
UGTR [yang2021uncertainty] ICCV 21 F 0.086 0.784 0.822 0.684 0.036 0.817 0.852 0.666 0.052 0.839 0.874 0.747
MGL-R [zhai2021mutual] CVPR 21 F 0.088 0.775 0.812 0.673 0.035 0.814 0.851 0.666 0.052 0.833 0.867 0.740
PFNet [mei2021camouflaged] CVPR 21 F 0.085 0.782 0.841 0.695 0.040 0.800 0.877 0.660 0.053 0.829 0.887 0.745
UJSC [li2021uncertainty] CVPR 21 F 0.073 0.800 0.859 0.728 0.035 0.809 0.884 0.684 0.047 0.842 0.898 0.771
BSA-Net [zhu2022can] AAAI 22 F 0.079 0.794 0.851 0.763 0.034 0.818 0.891 0.738 0.048 0.841 0.897 0.808
ZoomNet [pang2022zoom] CVPR 22 F 0.066 0.820 0.892 0.752 0.029 0.838 0.911 0.729 0.043 0.853 0.896 0.784
SINetv2 [fan2021concealed] TPAMI 22 F 0.070 0.820 0.895 0.782 0.037 0.815 0.906 0.718 0.048 0.847 0.914 0.805
HitNet [hu2023high] AAAI 23 F 0.055 0.849 0.906 0.831 0.023 0.871 0.935 0.823 0.037 0.875 0.926 0.853
SAM-Ada. [chen2023sam] ICCV 23 F 0.070 0.847 0.873 0.765 0.025 0.883 0.918 0.801 - - - -
FEDER [he2023camouflaged] CVPR 23 F 0.071 0.802 0.898 0.781 0.032 0.822 0.905 0.751 0.044 0.808 0.871 0.769
FSPNet [huang2023feature] CVPR 23 F 0.050 0.856 0.899 0.830 0.026 0.851 0.895 0.769 0.035 0.879 0.915 0.843
DINet [sun2024frequency] ECCV 24 F 0.068 0.821 0.874 0.755 0.031 0.832 0.903 0.729 0.043 0.856 0.909 0.794
CamoDiffusion [chen2024camodiffusion] AAAI 24 F 0.042 0.878 0.940 0.853 0.020 0.881 0.944 0.814 0.029 0.893 0.942 0.859
CamoFormer [yin2024camoformer] TPAMI 24 F 0.046 0.872 0.929 0.854 0.023 0.869 0.932 0.811 0.030 0.891 0.939 0.868
MCRNet [zhang2025mamba] IJCV 25 F 0.041 0.883 0.939 0.849 0.020 0.880 0.941 0.812 0.029 0.895 0.943 0.857
SCSOD [yu2021structure] AAAI 21 S 0.102 0.713 0.795 0.618 0.055 0.710 0.805 0.546 - - - -
CRNet [he2023weakly] AAAI 23 S 0.092 0.735 0.815 0.641 0.049 0.733 0.832 0.576 0.063 0.775 0.855 0.688
SAM [kirillov2023segment] ICCV 23 - 0.132 0.684 0.687 0.606 0.050 0.783 0.798 0.701 0.078 0.767 0.776 0.696
SAM-S [kirillov2023segment] ICCV 23 S 0.105 0.731 0.774 - 0.046 0.772 0.828 - 0.071 0.763 0.832 -
WS-SAM [hec2023weakly] NeurIPS 23 S 0.092 0.759 0.818 0.667 0.038 0.803 0.878 0.680 0.052 0.829 0.886 0.757
SAM-COD [chen2024sam] ECCV 24 S 0.060 0.836 0.903 0.779 0.029 0.833 0.904 0.728 0.039 0.859 0.912 0.795
@@ IV-A1 Implementation Details
We implement our model using PyTorch [paszke2019pytorch] and conduct all experiments on a single NVIDIA RTX 3090 GPU. In the first stage, we adopt SAM [kirillov2023segment] with a ViT-H backbone and Qwen2.5-VL-7B [Qwen2.5-VL] as the multimodal large language model (MLLM). In the second stage, following previous studies [huang2023feature, xie2023frequency], we employ the base version of ViT [dosovitskiy2020image] pretrained with the DeiT strategy [touvron2021training] as the transformer encoder.
For optimization, we use stochastic gradient descent (SGD) [ruder2016overview] with a momentum of 0.9 and a weight decay of $5\times 10^{-4}$. The batch size is set to 4, and the initial learning rate is 0.03. The learning rate is linearly warmed up during the first 20 epochs and subsequently decayed using a cosine annealing schedule. For loss configurations, the weight $w_{n}$ of the non-scribble class in $\mathcal{L}_{\text{Scrib}}$ is set to 0.02, and the modulation factor $\gamma$ in $\mathcal{L}_{\text{Debias}}$ is 0.9. In the overall loss formulation, the balancing coefficients are set to $\alpha=2$ and $\beta=0.5$.
@@ IV-A2 Datasets
To evaluate the effectiveness of our proposed method, we leverage the scribble-annotated S-COD dataset [he2023weakly], which is specifically designed for weakly-supervised camouflaged object detection, to generate pseudo masks and train our model. The trained model is then evaluated on three widely used benchmark datasets: CAMO [le2019anabranch], COD10K [fan2021concealed], and NC4K [lv2021simultaneously]. CAMO consists of 1,000 training images and 250 testing images. COD10K is a larger and more challenging dataset containing 3,040 training images and 2,026 testing images. NC4K serves as an additional benchmark, comprising 4,121 testing images without a training split.
@@ IV-A3 Evaluation Metrics
Following existing works [hec2023weakly, chen2024sam, he2023weakly], we adopt four commonly used metrics for evaluation: mean absolute error (MAE), structure measure ($S_{m}$) [fan2017structure], mean E-measure ($E_{m}$) [fan2021cognitive], and weighted F-measure ($F_{\beta}^{w}$) [margolin2014evaluate].
@@ IV-B1 Quantitative Comparison
To verify the effectiveness of our proposed framework, we compare it against state-of-the-art fully supervised and weakly-supervised methods, including SINet [fan2020camouflaged], UGTR [yang2021uncertainty], MGL-R [zhai2021mutual], PFNet [mei2021camouflaged], UJSC [li2021uncertainty], BSA-Net [zhu2022can], ZoomNet [pang2022zoom], SINetv2 [fan2021concealed], HitNet [hu2023high], SAM-Ada [chen2023sam], FEDER [he2023camouflaged], FSPNet [huang2023feature], DINet [sun2024frequency], CamoDiffusion [chen2024camodiffusion], CamoFormer [yin2024camoformer], MCRNet [zhang2025mamba], SCSOD [yu2021structure], CRNet [he2023weakly], SAM [kirillov2023segment], SAM-S [kirillov2023segment], WS-SAM [hec2023weakly], and SAM-COD [chen2024sam].
From the results in Table I, it is evident that ${D}^{3}$ETOR consistently outperforms all scribble-supervised baselines across all datasets. For example, on CAMO, it achieves an MAE of 0.045, $S_{m}$ of 0.868, $E_{m}$ of 0.928, and $F_{\beta}^{w}$ of 0.823, surpassing the second-best scribble-supervised method SAM-COD [chen2024sam] (MAE 0.060, $S_{m}$ 0.836, $E_{m}$ 0.903, $F_{\beta}^{w}$ 0.779) by a clear margin.
Furthermore, ${D}^{3}$ETOR demonstrates competitive or superior performance compared to most fully supervised approaches, only slightly behind CamoDiffusion [chen2024camodiffusion] and MCRNet [zhang2025mamba] on certain metrics. Overall, these results indicate that ${D}^{3}$ETOR establishes a new state-of-the-art for camouflaged object detection in weakly and even fully supervised scenarios using only scribble annotations.
@@ IV-B2 Qualitative Comparison
As illustrated in Figure 8, we select all existing weakly supervised approaches (i.e., CRNet [he2023weakly], WS-SAM [hec2023weakly], and SAM-COD [chen2024sam]) to visually compare against our proposed ${D}^{3}$ETOR. The qualitative results indicate that ${D}^{3}$ETOR accurately segments camouflaged objects under a variety of challenging conditions, including small-scale targets, intricate structural details, high visual resemblance to the background, fuzzy object boundaries, and complex surrounding contexts. In contrast, other methods fail to capture complete object regions or produce predictions with blurred details. These qualitative analyses provide an intuitive demonstration of the effectiveness and robustness of ${D}^{3}$ETOR.